OpenStudy (anonymous):

Help solving (x-8) (4x+1)=0

OpenStudy (anonymous):

(x-8) (4x+1)=0 x-8 = 0 x = 8 AND 4x+1 = 0 4x = -1 x = -1/4

OpenStudy (comm.dan):

What are the directions? @Virgo91684

OpenStudy (anonymous):

Thanks @tomo

OpenStudy (anonymous):

your welcome

OpenStudy (comm.dan):

@tomo is correct.